# From Hypothesis to Publication: A Comprehensive Survey of AI-Driven Research Support Systems

### v3

@@ 3 Hypothesis Validation
In scientific research, any proposed hypothesis must undergo rigorous validation to establish its validity. In some studies, this process is also referred to as ’falsification’ (Liu et al., 2024d; Huang et al., 2025). As illustrated in Figure  4, this section explores the application of AI in verifying scientific hypothesis through three methodological components: Scientific Claim Verification, Theorem Proving, and Experiment Validation.
@@ 3.1 Scientific Claim Verification
Scientific claim verification, also referred to as scientific fact-checking or scientific contradiction detection, aims to assess the veracity of claims related to scientific knowledge. This process assists scientists in verifying research hypothesis, discovering evidence, and advancing scientific work  (Wadden et al., 2020; Vladika and Matthes, 2023; Skarlinski et al., 2024). Research on scientific claim verification primarily focuses on three key elements: the claim, the evidence, and the validity of the claim  (Dmonte et al., 2024).
@@
Studies have highlighted that certain claims lack supporting evidence  (Wührl et al., 2024a), while others have demonstrated the ability to perform claim-evidence alignment without annotated data  (Bazaga et al., 2024). Additionally, methods such as HiSS  (Zhang and Gao, 2023) and ProToCo  (Zeng and Gao, 2023) proposed generating multiple claim variants to enhance verification.
@@
Existing research has explored various aspects related to evidence, including evidentiary sources  (Vladika and Matthes, 2024a), retrieval configurations  (Vladika and Matthes, 2024b), strategies for identifying and mitigating flawed evidence  (Glockner et al., 2022; Wührl et al., 2024b; Glockner et al., 2024a), and approaches for processing sentence-level  (Pan et al., 2023b) versus document-level indicators  (Wadden et al., 2022b).
@@
In the verification results generation phase, MAGIC (Kao and Yen, 2024) and SERIf (Cao et al., 2024b) proposed utilizing LLMs to synthesize evidence into more comprehensive information. FactKG  (Kim et al., 2023) and  Muharram and Purwarianti (2024) structured evidence into knowledge graphs, enabling claim attribution  (Dammu et al., 2024; Wu et al., 2023). Furthermore,  Atanasova et al. (2020); Krishna et al. (2022); Pan et al. (2023a); Eldifrawi et al. (2024); Zhang et al. (2024b) advocated for generating explanatory annotations alongside experimental outcomes during the verification process. Meanwhile,  Das et al. (2023); Altuncu et al. (2023) emphasized the critical role of domain expertise in ensuring accurate verification.
@@ 3.2 Theorem Proving
Theorem proving constitutes a subtask of logical reasoning, aimed at reinforcing the validity of the underlying theory within a hypothesis  (Pease et al., 2019; Yang et al., 2023c; Li et al., 2024e).
Following the proposal of GPT-f  (Polu and Sutskever, 2020) to utilize generative language models for theorem proving, researchers initially combined search algorithms with language models  (Lample et al., 2022; Wang et al., 2023b). However, a limitation in search-based approaches was later identified by  Wang et al. (2024a), who highlighted their tendency to explore insignificant intermediate conjectures. This led some teams to abandon search algorithms entirely. Subsequently, alternative methods emerged, such as the two-stage framework proposed by Jiang et al. (2023) and Lin et al. (2024), which prioritized informal conceptual generation before formal proof construction. Thor  (Jiang et al., 2022a) introduced theorem libraries to accelerate proof generation, an approach enhanced by Logo-power  (Wang et al., 2024b) through dynamic libraries. Studies like Baldur  (First et al., 2023), Mustard  (Huang et al., 2024c), and DeepSeek-Prover  (Xin et al., 2024) demonstrated improvements via targeted data synthesis and fine-tuning, though COPRA  (Thakur et al., 2024) questioned their generalizability and proposed an environment-agnostic alternative. Complementary strategies included theoretical decomposition into sub-goals  (Zhao et al., 2023b) and leveraging LLMs as collaborative assistants in interactive environments  (Song et al., 2024).
@@ 3.3 Experiment Validation
Experiment validation involves designing and conducting experiments based on the hypothesis. The empirical validity of the hypothesis is then determined through analysis of the experimental results (Huang et al., 2024b).
Experiment validation represents a time-consuming component of scientific research. Recent advancements in LLMs have enhanced their ability to plan and reason about experimental tasks  (Kambhampati et al., 2024), prompting researchers to use these models for designing and implementing experiments  (Ruan et al., 2024b). To ensure accuracy, studies such as  Zhang et al. (2023) and  Arlt et al. (2024) imposed input/output constraints, though this reduced generalizability. To address this,  Boiko et al. (2023); Bran et al. (2024); Huang et al. (2024a) integrated tools to expand model capabilities. Full automation was achieved by  Ni and Buehler (2023); Li et al. (2024a); Lu et al. (2024) through prompt-guided multi-agent collaboration.  Madaan et al. (2023); Yuan et al. (2025) further highlighted that the integration of feedback mechanisms demonstrated potential for enhancing design quality, while  Zhang et al. (2024a); Liu et al. (2024c); Ni et al. (2024) employed experimental outcomes to refine hyperparameter configurations, and  Szymanski et al. (2023); Li et al. (2024d); Baek et al. (2024) leveraged agent-generated analytical insights to facilitate iterative hypothesis refinement. In contrast, social science research often uses LLMs as experimental subjects to simulate human participants  (Liu et al., 2023b; Manning et al., 2024; Mou et al., 2024).
@@ 4.1 Manuscript Writing
Manuscript writing, also referred to as scientific or research writing. At this stage, researchers articulate the hypothesis they have formulated and the results they have validated in the form of a scholarly paper. This process is crucial, as it not only disseminates findings but also deepens researchers’ understanding of their work  (Colyar, 2009).
Early shared tasks focused on assisting researchers in writing or analyzing linguistic features (Dale and Kilgarriff, 2010; Daudaravicius, 2015). Recent advances have led to three main directions: citation text generation, related work generation, and complete manuscript generation.
@@
A subset of research on AI in scientific writing has focused on citation text generation, which addresses the academic need for referencing prior work while mitigating model inaccuracies  (Gao et al., 2023b; Gu and Hahnloser, 2023). For instance,  Wang et al. (2022b) developed an automated citation generation system by integrating manuscript content with citation graphs. However, its reliance on rigid template-based architectures led to inflexible citation formats. This limitation motivated subsequent studies to propose incorporating citation intent as a control parameter during text generation, aiming to improve contextual relevance and rhetorical adaptability  (Yu et al., 2022; Jung et al., 2022; Koo et al., 2023; Gu and Hahnloser, 2024).
@@
In contrast to citation text generation, several studies have focused on related work generation in scholarly writing, emphasizing the production of multiple citation texts and the systematic analysis of inter-citation relationships  (Li and Ouyang, 2022, 2024). The ScholaCite framework  (Martin-Boyle et al., 2024) leveraged GPT-4 to cluster citation sources and generate draft literature review sections, although it required manually provided reference lists. By contrast, the UR3WG system  (Shi et al., 2023) adopted a retrieval-augmented architecture integrated with large language models to autonomously acquire relevant references. To improve the quality of generated related work sections,  Yu et al. (2024b) utilized GNNs to model complex relational dynamics between target manuscripts and cited literature, while  Nishimura et al. (2024) initiative advocated for explicit novelty assertions regarding referenced publications.
@@
The aforementioned investigations primarily focused on specific components of scientific writing, while a study by  Lai et al. (2024) explored the progressive generation of complete manuscripts via structured workflows. The AI-Scientist system  (Lu et al., 2024) further introduced section-wise self-reflection mechanisms to enhance compositional coherence. Several studies emphasized human-AI collaborative frameworks for improving writing efficiency  (Lin, 2024; Feng et al., 2024; Ifargan et al., 2024), whereas  Tang et al. (2024a) concentrated on enabling personalized content generation in multi-author collaborative environments. Following initial manuscript drafting, subsequent text revision processes were systematically examined  (Du et al., 2022b; Jourdan et al., 2023; Dang et al., 2025). The OREO system  (Li et al., 2022) utilized attribute classification for iterative in-situ editing, while  Du et al. (2022a); Pividori and Greene (2024) incorporated researcher feedback loops for progressive text optimization. Notably,  Kim et al. (2022); Chamoun et al. (2024); D’Arcy et al. (2024b) proposed replacing manual feedback with automated evaluation metrics.
@@ 4.2 Peer Review
Peer review serves as a critical mechanism for improving the quality of academic manuscripts through feedback and evaluation, forming the cornerstone of quality control in scientific research. However, the process is hindered by its slow pace, high time consumption, and increasing strain due to the growing academic workload  (Lin et al., 2023a; Kousha and Thelwall, 2024; Thelwall and Yaghi, 2024). To address these challenges and enhance manuscript quality, researchers have investigated the application of AI in peer review  (Yuan et al., 2022; Liu and Shah, 2023; Niu et al., 2023; Kuznetsov et al., 2024; Thakkar et al., 2025). Peer review can be categorized into two main types: paper review generation and meta-review generation.
@@
In paper review generation, reviewers provide both scores and evaluations for manuscripts. For instance,  Setio and Tsuchiya (2022) formulated score prediction as a regression task,  Muangkammuen et al. (2022) utilized semi-supervised learning, and  Couto et al. (2024) treated the task as a classification problem to evaluate the alignment between manuscripts and review criteria. While these approaches focused on label prediction for paper reviews,  Yuan and Liu (2022) extended the scope by directly generating reviews through the construction of a concept graph integrated with a citation graph.
Subsequently, a pilot study conducted by  Robertson (2023) demonstrated the capability of GPT-4 to generate paper reviews. Further investigations, such as those by AI-Scientist  (Lu et al., 2024) and  Liang et al. (2023), evaluated its performance as a review agent. Additionally, systems like MARG  (D’Arcy et al., 2024a) and SWIF2T  (Chamoun et al., 2024) employed multi-agent frameworks to generate reviews via internal discussions and task decomposition. In contrast, AgentReview  (Jin et al., 2024) and  Tan et al. (2024) modeled the review process as a dynamic, multi-turn dialogue. Furthermore, CycleResearcher  (Weng et al., 2024) and OpenReviewer  (Idahl and Ahmadi, 2024) fine-tuned models for comparative reviews and structured outputs aligned with conference guidelines.
@@
In meta-review generation, chairs are tasked with identifying a paper’s core contributions, strengths, and weaknesses while synthesizing expert opinions on manuscript quality. Meta-reviews are conceptualized as abstractions of comments, discussions, and paper abstracts  (Li et al., 2023).  Santu et al. (2024) investigated the use of LLMs for automated meta-review generation, while  Zeng et al. (2023) proposed a guided, iterative prompting approach. MetaWriter  (Sun et al., 2024) utilized LLMs to extract key reviewer arguments, whereas GLIMPSE  (Darrin et al., 2024) and  Kumar et al. (2023) focused on reconciling conflicting statements to ensure fairness. Additionally,  Li et al. (2024b) introduced a three-layer sentiment consolidation framework for meta-review generation, and PeerArg  (Sukpanichnant et al., 2024) integrated LLMs with knowledge representation to address subjectivity and bias via a multiparty argumentation framework (MPAF). DeepReview (Zhu et al., 2025) generates a comprehensive meta-review by simulating expert evaluation across multiple dimensions.

### v4

@@ 3 Hypothesis Validation
In scientific research, any proposed hypothesis must undergo rigorous validation to establish its validity. In some studies, this process is also referred to as ’falsification’ (2411ZijunLiuAIGS; 2502KexinHuangAutomated). As illustrated in Figure  4, this section explores the application of AI in verifying scientific hypothesis through three methodological components: Scientific Claim Verification, Theorem Proving, and Experiment Validation.
@@ 3.1 Scientific Claim Verification
Scientific claim verification, also referred to as scientific fact-checking or scientific contradiction detection, aims to assess the veracity of claims related to scientific knowledge. This process assists scientists in verifying research hypothesis, discovering evidence, and advancing scientific work  (2004DavidWaddenFact; 2305JurajVladikaScientific; 2409MichaelDSkarlinskiLanguage). Research on scientific claim verification primarily focuses on three key elements: the claim, the evidence, and the validity of the claim  (2408AlphaeusDmonteClaim).
@@
Studies have highlighted that certain claims lack supporting evidence  (2402AmelieWhrlWhat), while others have demonstrated the ability to perform claim-evidence alignment without annotated data  (2309AdrinBazagaUnsupervised). Additionally, methods such as HiSS  (2310XuanZhangTowards) and ProToCo  (2306FengzhuZengPrompt) proposed generating multiple claim variants to enhance verification.
@@
Existing research has explored various aspects related to evidence, including evidentiary sources  (2402JurajVladikaComparing), retrieval configurations  (2404JurajVladikaImproving), strategies for identifying and mitigating flawed evidence  (2210MaxGlocknerMissing; 2402AmelieWhrlUnderstanding; 2408MaxGlocknerGrounding), and approaches for processing sentence-level  (2309LiangmingPanInvestigating) versus document-level indicators  (2112DavidWaddenMultiVerS).
@@
In the verification results generation phase, MAGIC (24WeiYuKaoMAGIC) and SERIf (2402YupengCaoCan) proposed utilizing LLMs to synthesize evidence into more comprehensive information. FactKG  (2305JihoKimFactKG) and  2409AriefPurnamaMuharramEnhancing structured evidence into knowledge graphs, enabling claim attribution  (2403PreetamPrabhuSrikarDammuClaimVer; 23JinxuanWuCharacterizing). Furthermore,  2004PepaAtanasovaGenerating; 2108AmrithKrishnaProoFVer; 2305LiangmingPanFact-Checking; 2407IslamEldifrawiAutomated; 2410XiaochengZhangAugmenting advocated for generating explanatory annotations alongside experimental outcomes during the verification process. Meanwhile,  2301AnubrataDasThe; 2305EnesAltuncuaedFaCT emphasized the critical role of domain expertise in ensuring accurate verification.
@@ 3.2 Theorem Proving
Theorem proving constitutes a subtask of logical reasoning, aimed at reinforcing the validity of the underlying theory within a hypothesis  (pease2019importance; 2303ZonglinYangLogical; 2404ZhaoyuLiA).
Following the proposal of GPT-f  (2009StanislasPoluGenerative) to utilize generative language models for theorem proving, researchers initially combined search algorithms with language models  (2205GuillaumeLampleHyperTree; 23DT-SolverHaimingWang). However, a limitation in search-based approaches was later identified by  2405HaimingWangProving, who highlighted their tendency to explore insignificant intermediate conjectures. This led some teams to abandon search algorithms entirely. Subsequently, alternative methods emerged, such as the two-stage framework proposed by 2210AlbertQiaochuJiangThe and 2407HaohanLinLean-STaR, which prioritized informal conceptual generation before formal proof construction. Thor  (2407AlbertQiaochuJiangAdvances) introduced theorem libraries to accelerate proof generation, an approach enhanced by Logo-power  (2310HaimingWangLEGO-Prover) through dynamic libraries. Studies like Baldur  (2303EmilyFirstBaldur), Mustard  (2303YinyaHuangMUSTARD), and DeepSeek-Prover  (2410HuajianXinDeepSeek-Prover) demonstrated improvements via targeted data synthesis and fine-tuning, though COPRA  (24ThakurAn) questioned their generalizability and proposed an environment-agnostic alternative. Complementary strategies included theoretical decomposition into sub-goals  (2305XueliangZhaoDecomposing) and leveraging LLMs as collaborative assistants in interactive environments  (2404PeiyangSongTowards).
@@ 3.3 Experiment Validation
Experiment validation involves designing and conducting experiments based on the hypothesis. The empirical validity of the hypothesis is then determined through analysis of the experimental results (2310QianHuangMLAgentBench).
Experiment validation represents a time-consuming component of scientific research. Recent advancements in LLMs have enhanced their ability to plan and reason about experimental tasks  (2402SubbaraoKambhampatiPosition), prompting researchers to use these models for designing and implementing experiments  (24RuanYixiangAn). To ensure accuracy, studies such as  2305ShujianZhangAutoMLGPT and  2406SrenArltMetaDesigning imposed input/output constraints, though this reduced generalizability. To address this,  23DaniilABoikoAutonomous; 2304AndresMBranAugmenting; 2404KaixuanHuangCRISPRGPT integrated tools to expand model capabilities. Full automation was achieved by  2311BoNiMechAgents; 2410LongLiChain; 2408ChrisLuThe through prompt-guided multi-agent collaboration.  2303AmanMadaanSelfRefine; 2501YuanJiakangDolphin further highlighted that the integration of feedback mechanisms demonstrated potential for enhancing design quality, while  2304LeiZhangMLCopilot; 2402SiyiLiuLarge; 2411ZiqiNiMatPilot employed experimental outcomes to refine hyperparameter configurations, and  23SzymanskiNathanJautonomousAn; 2408RuochenLiMLRCopilot; 2404JinheonBaekResearchAgent leveraged agent-generated analytical insights to facilitate iterative hypothesis refinement. In contrast, social science research often uses LLMs as experimental subjects to simulate human participants  (2305RuiboLiuTraining; 2404ManningBenjaminSAutomated; 2412XinyiMouFrom).
@@ 4.1 Manuscript Writing
Manuscript writing, also referred to as scientific or research writing. At this stage, researchers articulate the hypothesis they have formulated and the results they have validated in the form of a scholarly paper. This process is crucial, as it not only disseminates findings but also deepens researchers’ understanding of their work  (09ColyarBecoming).
Early shared tasks focused on assisting researchers in writing or analyzing linguistic features (dale2010helping; daudaravicius2015automated). Recent advances have led to three main directions: citation text generation, related work generation, and complete manuscript generation.
@@
A subset of research on AI in scientific writing has focused on citation text generation, which addresses the academic need for referencing prior work while mitigating model inaccuracies  (2305TianyuGaoEnabling; 2306NianlongGuSciLit). For instance,  2206YifanWangDisenCite developed an automated citation generation system by integrating manuscript content with citation graphs. However, its reliance on rigid template-based architectures led to inflexible citation formats. This limitation motivated subsequent studies to propose incorporating citation intent as a control parameter during text generation, aiming to improve contextual relevance and rhetorical adaptability  (22YuMengxiaScientific; 22JungShingYunIntent; 2304RyanKooDecoding; 24GuNianlongControllable).
@@
In contrast to citation text generation, several studies have focused on related work generation in scholarly writing, emphasizing the production of multiple citation texts and the systematic analysis of inter-citation relationships  (2201XiangciLiAutomatic; 2404XiangciLiRelated). The ScholaCite framework  (2402MartinBoyleAnnaShallow) leveraged GPT-4 to cluster citation sources and generate draft literature review sections, although it required manually provided reference lists. By contrast, the UR3WG system  (23ZhengliangShiTowards) adopted a retrieval-augmented architecture integrated with large language models to autonomously acquire relevant references. To improve the quality of generated related work sections,  24LuyaoYuReinforced utilized GNNs to model complex relational dynamics between target manuscripts and cited literature, while  24NishimuraToward initiative advocated for explicit novelty assertions regarding referenced publications.
@@
The aforementioned investigations primarily focused on specific components of scientific writing, while a study by  2408YuxuanLaiInstruct explored the progressive generation of complete manuscripts via structured workflows. The AI-Scientist system  (2408ChrisLuThe) further introduced section-wise self-reflection mechanisms to enhance compositional coherence. Several studies emphasized human-AI collaborative frameworks for improving writing efficiency  (2310LinZhichengTechniques; 2412KJKevinFengCocoa; 2404TalIfarganAutonomous), whereas  2406XiangruTangStep concentrated on enabling personalized content generation in multi-author collaborative environments. Following initial manuscript drafting, subsequent text revision processes were systematically examined  (DuR0KLK22; 2303LaneJourdanText; DangSBG25). The OREO system  (2204JingjingLiText) utilized attribute classification for iterative in-situ editing, while  2204WanyuDuRead; 24MiltonPividoriA incorporated researcher feedback loops for progressive text optimization. Notably,  KimDR0K22; 2405EricChamounAutomated; 2306MikeDArcyARIES proposed replacing manual feedback with automated evaluation metrics.
@@ 4.2 Peer Review
Peer review serves as a critical mechanism for improving the quality of academic manuscripts through feedback and evaluation, forming the cornerstone of quality control in scientific research. However, the process is hindered by its slow pace, high time consumption, and increasing strain due to the growing academic workload  (2111JialiangLinAutomated; 23KayvanKoushaArtificial; 2411MikeThelwallEvaluating). To address these challenges and enhance manuscript quality, researchers have investigated the application of AI in peer review  (22WeizheYuanCan; 2306RyanLiuReviewerGPT; 2309LiangNiuUnveiling; 2405IliaKuznetsovWhat; thakkar2025can). Peer review can be categorized into two main types: paper review generation and meta-review generation.
@@
In paper review generation, reviewers provide both scores and evaluations for manuscripts. For instance,  22BasukiSetioThe formulated score prediction as a regression task,  22PanitanMuangkammuenExploiting utilized semi-supervised learning, and  2406PauloHenriqueCoutoRelevAI treated the task as a classification problem to evaluate the alignment between manuscripts and review criteria. While these approaches focused on label prediction for paper reviews,  22WeizheYuanKID extended the scope by directly generating reviews through the construction of a concept graph integrated with a citation graph.
Subsequently, a pilot study conducted by  2307ZacharyRobertsonGPT4 demonstrated the capability of GPT-4 to generate paper reviews. Further investigations, such as those by AI-Scientist  (2408ChrisLuThe) and  2310WeixinLiangCan, evaluated its performance as a review agent. Additionally, systems like MARG  (2401MikeDArcyMARG) and SWIF2T  (2405EricChamounAutomated) employed multi-agent frameworks to generate reviews via internal discussions and task decomposition. In contrast, AgentReview  (2406YiqiaoJinAgentReview) and  2406ChengTanPeer modeled the review process as a dynamic, multi-turn dialogue. Furthermore, CycleResearcher  (2411YixuanWengCycleResearcher) and OpenReviewer  (2412MaximilianIdahlOpenReviewer) fine-tuned models for comparative reviews and structured outputs aligned with conference guidelines.
@@
In meta-review generation, chairs are tasked with identifying a paper’s core contributions, strengths, and weaknesses while synthesizing expert opinions on manuscript quality. Meta-reviews are conceptualized as abstractions of comments, discussions, and paper abstracts  (2305MiaoLiSummarizing).  2402ShubhraKantiKarmakerSantuPrompting investigated the use of LLMs for automated meta-review generation, while  2305QiZengMeta proposed a guided, iterative prompting approach. MetaWriter  (24LuSunMetaWriter) utilized LLMs to extract key reviewer arguments, whereas GLIMPSE  (2406MaximeDarrinGLIMPSE) and  2310SandeepKumarWhen focused on reconciling conflicting statements to ensure fairness. Additionally,  2402MiaoLiA introduced a three-layer sentiment consolidation framework for meta-review generation, and PeerArg  (2409PurinSukpanichnantPeerArg) integrated LLMs with knowledge representation to address subjectivity and bias via a multiparty argumentation framework (MPAF). DeepReview (zhu2025deepreview) generates a comprehensive meta-review by simulating expert evaluation across multiple dimensions.
@@ 5 Benchmarks
Task Benchmark Domain Size Input Output Metric
SCHOLAT 20WeishengLiAcademic Research Paper Recommendation 34,518 - - Recall, Precission, F1-score
ACL selection network tao2020paper Research Paper Recommendation 18,718 Topics Related Papers Accuracy
CiteSeer 21Kanghybrid Research Paper Recommendation 1,100 Paper Related Papers Correlation Coefficient
SciReviewGen 23TetsuKasanishiSciReviewGen Systematic Literature Review 10,000+ Abstracts literature review ROUGE
FacetSum 21RuiMengBringing Systematic Literature Review 60,024 Source Text+Facet Summary of Facet ROUGE
BigSurvey 22ShuaiqiLiuGenerating Systematic Literature Review 7,000+ Abstracts Survey Paragraph ROUGE, F1-score
SCHOLARQABENCH 2411AkariAsaiOpenScholar Systematic Literature Review 2,200 Question Answer with Citations
Accuracy, Coverage, Citations
+ Relevance, Usefulness
HiCaD 2304KunZhuHierarchical Systematic Literature Review 7,600 Reference Papers Catalogues
Catalogue Edit Distance Similarity (CEDS)
+ Catalogue Quality Estimate (CQE)
CLUSTREC-COVID 2408UriKatzKnowledge Systematic Literature Review 2,284 Titles, Abstracts Topic Clusters per Topic
CHIME 2407ChaoChunHsuCHIME Systematic Literature Review 2,174 Topic Hierarchies F1-score
24DerekFWongOverview Systematic Literature Review 700 Subject, Reference Title,Content -
MASSW 2406XingjianZhangMASSW Hypothesis Generation 152000 Context of Literature Hypothesis
BLEU, ROUGE, BERTScore,
+ Cosine Similarity, BLEURT
IdeaBench 2411SikunGuoIdeaBench Hypothesis Generation 2,374 Instruction, Background Information Hypothesis
Insight Score, BERTScore, Novelty,
+ LLM Similarity Rating, Feasibility
SCIMON 2305QingyunWangSciMON Hypothesis Generation - Background Context Idea
ROUGE, BERTScore
+BARTScore, Novelty
MOOSE2309ZonglinYangLarge Hypothesis Generation 50 Background, Inspiration Hypothesis
Validness, Novelty
+ Helpfulness
DISCOVERYBENCH 2407DiscoveryBenchBodhisattwa Hypothesis Generation 1,167 Data Discovery
Hypothesis Match Score
Hypothesis LiveIdeaBench 2412LiveIdeaBench Hypothesis Generation - Scientific Keywords Idea
Originality, Feasibility
+ Fluency, Flexibilit
Formulation 2409SandeepKumarCan Hypothesis Generation 100 Paper without Future Work Idea
Idea Alignment Score, Idea Distinctness Index
SciRIFF 2406DavidWaddenSciRIFF Scientific Claim Verification 137,000 Evidence, Task prompt Structured Paragraph F1, BLEU
SCIFACT 2004DavidWaddenFact Scientific Claim Verification 1,409 Claim, Evidence Rationale Sentences, Label Precision, Recall, Micro-F1
SCIFACT-OPEN WaddenLKCBWH22 Scientific Claim Verification 279 Claim, Evidence Rationale Sentences, Label Precision, Recall,Micro-F1
MISSCI Glockner0NG24 Scientific Claim Verification 435 Claim, Premise, Context Verification
Micro F1-score,P@1,Arg@1
+ METEOR Score,BERTScore
+NLI-A, NLI-S, Matches@1
FEVER ThorneVCM18 Scientific Claim Verification 185,445 Claim, Evidence Label, Necessary Evidence
F1-Score,Oracle Accuracy
+ Accuracy,Recall
XClaimCheck 24WeiYuKaoMAGIC Scientific Claim Verification 16,177 Claim, Evidence Label, Argument Macro-F1, Accuracy
HEALTHVER SarroutiAMD21 Scientific Claim Verification 14330 Claim, Evidence Label
Macro Precision, Macro Recall
+ Macro F1-score, Accuracy
QuanTemp VAAS24 Scientific Claim Verification 15,514 Claim, Evidence Label
Weighted-F1 Score, Macro-F1, BLEU,
+ BERTScore, Cohen’s Kappa Score
+ Human Evaluation
SCITAB LuPLNK23 Scientific Claim Verification 1,225 Claim, Evidence Label
Macro-F1
Check-COVID WangHCASM23 Scientific Claim Verification 1,504 Claim Evidence
Accuracy, Precision, Recall, Macro-F1
HealthFC VladikaSM24 Scientific Claim Verification 750 Claim, Evidence Label
Precision, Recall, F1-Macro
FACTKG 2305JihoKimFactKG Scientific Claim Verification 108,000 Claim, Evidence Label
Accuracy
BEAR-FACT 2402AmelieWhrlWhat Scientific Claim Verification 1,448
Claim, Evidence
+Entity/Relation Information
Label
F1-Score
MINIF2F ZhengHP22 Theorem Proving 488 Problem, Theorem Proof
Pass Rate
FIMO 2309ChengwuLiuFIMO Theorem Proving 149 Problem, Theorem, statements Proof
Pass Rate
LeanDojo YangSGCSYGPA23 Theorem Proving 98,734 Problem, Theorem Proof
R@k, MRR, Pass Rate
Lean-github 2407LEANGitHub Theorem Proving 28,597 Problem, Theorem Proof
Accuracy, Pass Rate
TRIGO-real XiongSYWYLLGCHZ23 Theorem Proving 427 Problem, Theorem Proof
Pass Rate, Accuracy, EM@n
TRIGO-web XiongSYWYLLGCHZ23 Theorem Proving 453 Problem, Theorem Proof
Pass Rate, Accuracy, EM@n
TRIGO-gen XiongSYWYLLGCHZ23 Theorem Proving - Problem, Theorem Proof
Pass Rate, Accuracy, EM@n
CoqGym YangD19 Theorem Proving 71,000 Problem, Theorem Proof
Success Rate
MLAgentBench 2310QianHuangMLAgentBench Experiment Validation 13 - -
Competence, Efficiency
AAAR-1.0 2410RenzeLouAAAR10 Experiment Validation - Instance, Papers Design, Explanation
S-F1, S-Precision, S-Recall
+ S-Match, ROUGE
TASKBENCH 0001STZRY00Z24 Experiment Validation 17,331 - -
ROUGE, t-F1, v-F1
+Normalized Edit Distance
Spider2-V CaoLWCFGXZHMXXZ24 Experiment Validation 494 Task Experiment Execution
Success Rate
CORE-Bench 2409ZacharySSiegelCOREBench Experiment Validation 270 Task Requirements Experiment Result
Accuracy
LAB-Bench 2407LABBench Experiment Validation 2400 Multiple-choice Question Answer
Accuracy, Precision, Coverage
PaperBench starace2025paperbench Experiment Validation 20 Paper, Additional Information Code
Replication Score
Hypothesis SUPER BoginYG0BCSK24 Experiment Validation 801 Task Requirements -
Accuracy, Landmark-Based Evaluation
Validation ScienceAgentBench ChenCNZWYLLWLDX25 Experiment Validation 102
Task Instruction, Dataset Information
+Expert-Provided Knowledge
Program
Valid Execution Rate, Success Rate, CodeBERTScore, API Cost
SciCap+ 23ZhishenYangSciCap Manuscript Writing 414,000
Figure, OCR tokens
+ Mention Paragraph
Caption
BLEU, ROUGE, METEOR
+ CIDEr, SPICE
AAN Corpus RadevMQA13 Manuscript Writing - - - -
SciSummNet YasunagaKZFLFR19 Manuscript Writing 1,000 Paper,Citation Sentence Summary
ROUGE
CiteBench FunkquistK0G23 Manuscript Writing 358,765 Cited Papers, Context Citation Text
ROUGE, BERTScore
ALCE 2305TianyuGaoEnabling Manuscript Writing 3,000 Question Answer with Citations
Recall, Precision
GCite 2206YifanWangDisenCite Manuscript Writing 2,500 Citing/Cited Paper Citation Text
BLEU, ROUGE
ARXIVEDITS JiangXS22 Manuscript Writing 1,000 Sentence Pairs Sentence, Intent
Precision,Recall,F1-score
CASIMIR JourdanBHD24 Manuscript Writing 15,646 Original Sentence Revised Sentence
Exact-match (EM),SARI, BLEU,
+ ROUGE-L,Bertscore
ParaRev jourdan2025pararev Manuscript Writing 48,203 Original Paragraph Revised Paragraph
ROUGE-L,SARI
+ BertScore
SCHOLAWRITE 2502LingheWangScholaWrite Manuscript Writing 62,000 Before-text Writing Intention, After-text
F1-score, Lexical Diversity, Topic Consistency, Intention Coverage
MReD ShenCZBYS22 Peer Review 7,089 Reviews Meta-Review
ROUGE
ORSUMzeng2024scientific Peer Review 15,062 Reviews Meta-Review
ROUGE-L, BERTScore, FACTCC
+ SummaC, DiscoScore
PeerRead v1 KangADZKHS18 Peer Review 107,000 Reviews Accept/Reject
Accuracy
NLPeer DyckeKG23 Peer Review 5,000 Reviews,Paper
Review Score, Connection,
+ Review Category
MRSE, F1-macro
+ Precision, Recall
AMPERE HuaNBW19 Peer Review 400 Review Review with Type
Precision, Recall, F1-score
MOPRD LinSZCS23 Peer Review 6,578 Reviews,Paper
Editorial Decision, Review,
+ Meta-Review, Author Rebuttal
ROUGE, BARTScore
ARIES 2306MikeDArcyARIES Peer Review 1,720 Review Comment, Edits Comment-Edit Pairs
Precision, Recall, F1-score
ASAP-Review 22WeizheYuanCan Peer Review - Paper Review
Aspect Coverage, Aspect Recall,
+Semantic Equivalence
+Human: Recommendation Accuracy(RAcc),
+Informativeness(Info),Aspect-level,
+Constructiveness(ACon) and Summary accuracy
ReviewMT 2406ChengTanPeer Peer Review 26,841 Paper Review Dialogue
ROUGE,BLEU,METEOR
Manuscript ReAct ChoudharyMM21 Peer Review 6,250 Review Classification of Review
Accuracy
Publication PEERSUM 2305MiaoLiSummarizing Peer Review - Reviews Meta-Review
ROUGE,BERTScore,UniEval,ACC